OpenStudy (anonymous):

what is the sum of 7/9 + 1/6 + 11/12=

OpenStudy (anonymous):

You have to get the LCD of the denominators: 9=3^2; 6=2*3;12=2^2*3 Take the product of each distinct factor at its highest power, so\[LCD=2^2*3^2=36\]Then you have to multiply each fraction by the appropriate value to make the given denominator the LCD; of course you have to multiply the numerator by the same value as the denominator:\[\frac{7}{9}*\frac{4}{4}+\frac{1}{6}*\frac{6}{6}+\frac{11}{12}*\frac{3}{3}\]Then add the numerators\[\frac{28+6+33}{36}=\frac{67}{36}\]